OpenStudy (anonymous):

Solve for x. -6x + 11 -9x + 2

OpenStudy (solomonzelman):

\[-6x + 11 = -9x + 2 \]this?

OpenStudy (anonymous):

yes @solomonzelman

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

thx

OpenStudy (solomonzelman):

\[-6x+11\color{red} {+9x-11}=-9x+2\color{red} {+9x-11}\]\[3x=-9\]\[\color{blue} {x=?}\]